# Holey sheets: Double-Threshold Rupture of Draining Liquid Films

### v2

@@ Abstract
In this Letter, we show that an air bubble entrained in a draining liquid sheet can nucleate a hole long before molecular-scale forces become relevant. Axisymmetric direct numerical simulations performed with the Navier-Stokes and combined level-set-volume-of-fluid solver Basilisk C [25, 26] isolate this impurity-driven pathway under controlled conditions while retaining realistic rim dynamics. To focus on the essential physics, we replace the inflating bag geometry of air-blast breakup by an initially flat sheet of thickness $h_{0}$ that drains radially with $u(r)=\omega r$ (fig. 1b), a template that captures both external-airflow forcing in bag breakup [38] and inertia-driven drainage after drop impact [39]. The dynamics are governed by the Ohnesorge number $Oh=\eta/\sqrt{\rho\gamma R_{0}}$, the Bond number $Bo=\rho\omega^{2}R_{0}^{3}/\gamma$, and the non-dimensionlized offset $\chi/R_{0}$ of the bubble’s center from the centerlines of the flow. Here $\eta$ and $\rho$ are the liquid viscosity and density, $\gamma$ is the surface tension, and $R_{0}$ the bubble radius. The driving acceleration $\omega^{2}R_{0}$ acts as an effective radial gravity. In addition, the initial cavity distortion is characterized by an open angle $\theta$ (fig. 1c), see supplementary material [26] for details on drainage mechanism leading to this geometry.
@@
Our two key results are as follows: First, the entrained bubble grows a through-cavity that overcomes surface tension and nucleates a hole at micrometer-scale thicknesses, explaining rupture without invoking nanoscale physics. Second, the post-nucleation fate is deterministic and requires crossing a double-threshold for irreversible opening: the outward drainage must exceed a critical $Bo_{c}(Oh)$ and the initial cavity distortion – characterized by an opening angle $\theta$ – must exceed a critical $\theta_{c}(Oh)$. Otherwise, rims collide and the sheet heals. The opening-healing boundary collapses onto simple asymptotic scalings: $Bo_{c}=\mathcal{O}(1)$ for $Oh\ll 1$ and $Bo_{c}\sim Oh^{-2}$ for $Oh\gg 1$. These results rationalize the observed perforation of micron-thick sheets by internal defects and provide a predictive framework for impurity-induced breakup in natural and engineered fragmentation flows.
@@
Phenomenology – We first examine sheet-centered bubbles ($\chi/R_{0}=0$) across the $(Oh,Bo)$ space. As the sheet thins, its top and bottom interfaces meet at the bubble poles and a through-cavity opens. Capillary waves invert the cavity from convex to concave, after which two outcomes are observed (fig. 2a,b): either the rims retract and the hole expands, breaking the sheet, or capillarity drives rim collision and the sheet heals. A systematic sweep in $(Oh,Bo)$ reveals a sharp transition curve (fig. 2c): strong radial driving (large $Bo$) yields opening, whereas weaker driving permits healing unless $Bo$ exceeds a viscosity-dependent threshold. The location and slope of this boundary reflect the competition between capillary-driven rim closure and drainage-driven advection, anticipating the timescale analysis that follows. We additionally show that delayed nucleation–representing chemical/thermal heterogeneity–can be encoded by a larger initial cavity characterized by an opening angle $\theta$, which shifts the transition and thereby introduces geometry as a second way of control.
@@
Figure 2: As the liquid sheet drains radially and thins axially, the interfaces of the bubble and sheet merge to create a cavity. The time instants are shown for $Oh=0.1$, (a) and at small $Bo=10^{-3}$ where the sheet heals, while (b) at larger $Bo=10^{-2}$, the sheet opens up. The left panel depicts velocity magnitude $\|\bm{v}\|/V_{0}$, where $V_{0}=\sqrt{Bo\gamma/\rho R_{0}}$, and the black arrows depict the velocity direction. The right panel illustrates viscous dissipation $\xi=2Bo\left(\bm{\mathcal{D}:\mathcal{D}}\right)$, where $\bm{\mathcal{D}}=\left(\bm{\nabla u+\nabla u}^{T}\right)/2$ is the symmetric part of velocity gradient tensor. (c) The regime map in the log-log parameter space of OhBoOh-Bo. The transition lines at small OhOh are shown by a constant BoBo line in gray, while, at large OhOh, the transition is indicated by a black line with scaling $Oh\sim Bo^{-1/2}$.
@@
What happens in the limiting case of no external driving ($Bo=0$)? – In the absence of radial forcing ($Bo=0$), the liquid sheet inevitably heals for any finite $Oh$ – even extremely viscous sheets (large $Oh$) will eventually reseal, albeit very slowly as viscosity prolongs the rim collision time. This follows from our thresholds: as $Bo\to 0$, achieving opening would require $Oh\sim Bo^{-1/2}\to\infty$, so any finite $Oh$ lies in the healing regime. Our simulations confirm this universal healing at $Bo=0$, consistent with the classic geometric energy criterion of Taylor & Michael [40]. Using a toroidal model for the hole’s rim [41], they predicted that without external driving a circular hole can continue to expand only if its radius exceeds a purely geometric threshold. In particular, when the hole’s outer radius $R$ satisfies $R/h_{0}>\pi/4$ (so that the inner radius $R_{h}=R-h_{0}/2$ is about $0.29h_{0}$), the excess surface energy $\Delta\Xi$ becomes positive, signaling that the hole will grow rather than close. Holes smaller than this critical size have $\Delta\Xi<0$ and will heal. Notably, this threshold is independent of viscosity. Thus, at $Bo=0$ no finite-$Oh$ case can open unless the initial hole is above the critical $R/h_{0}$ – indeed, the cavities in our simulations were below this size, so they all heal.
In more realistic situations, additional physics can delay the initial rupture (for example, due to chemical or thermal heterogeneity), allowing the sheet to drain longer and thin more broadly before the hole forms. We represent such delays in our model by a larger initial cavity distortion characterized by an increased polar angle $\theta$ (fig. 1c). Simulations at $Bo=0$ with varied $\theta$ reveal a clear geometry-controlled threshold: for sufficiently large distortions (large $\theta$), the hole opens even without any driving, whereas for smaller initial distortions the hole heals (fig. 3b). This confirms the second threshold of the double-threshold framework: a sufficiently large initial cavity alone can trigger sheet rupture in the absence of external forcing. Moreover, at very high viscosity (large $Oh$), the opening-healing transition occurs at a nearly constant $\theta\approx 0.09\pi$, consistent with the $R_{h}/h_{0}\approx 0.29$ criterion seen in our simulations at large $Oh$. As $Oh$ decreases (inertia-dominated regime), this critical distortion $\theta_{c}$ rises, since added inertia aids rim closure—requiring a larger initial opening to overcome healing. Remarkably, sheets that do open exhibit non-monotonic cavity growth (fig. 3c). Surface tension first drives the cavity edges apart, causing radial cavity growth, but the induced inertia pulls the rim inward and temporarily reduces the cavity radius. However, the cavity remains large enough to overcome healing by this recoil and eventually opens with the Taylor–Culick mechanism.
@@
Figure 4: The draining liquid sheet of the bubble is placed asymmetrically with $\chi/R_{0}=0.1$. Time evolution has been shown for cases with $Oh=1$, (a) $Bo=10^{-4}$, and (b) $Bo=10^{-2}$. In the former case, the liquid bridge at the south pole replenishes and grows into a flat-shaped sheet due to capillary action, and then the sheet keeps on thinning due to radial drainage. Meanwhile, in the latter case, thinning due to radial drainage dominates before the bridge can be replenished by the damped capillary waves at large $Oh=1$. The left panel depicts velocity magnitude $\|\bm{v}\|/V_{0}$, where $V_{0}=\sqrt{Bo\gamma/\rho R_{0}}$, the black arrows depict the velocity direction. The right panel depicts viscous dissipation $\xi=2Bo\left(\bm{\mathcal{D}:\mathcal{D}}\right)$, where $\bm{\mathcal{D}}=\left(\bm{\nabla u+\nabla u}^{T}\right)/2$ is the symmetric part of velocity gradient tensor. (c) Sheet breakup time (normalized by the no-bubble case rupture time $\tau^{*}\sim\tau_{\gamma}/\sqrt{Bo}$) in the OhBoOh-Bo parameter space for $\chi/R_{0}=0.1$. The gray line depicts the transition line observed for symmetric cases ($\chi/R_{0}=0$), as shown in fig. 2.
@@
A trapped bubble – acting as a hydrophobic defect – thus rationalizes why rupture occurs far above molecular scales and provides predictive control knobs, namely driving strength and defect geometry, relevant to bag-breakup sprays and respiratory films. Future studies could test whether analogous thresholds govern rupture triggered by solid particles, oil droplets, or Marangoni-driven inhomogeneities. Chemical or thermal gradients may couple nonlinearly with radial drainage and shift $Bo_{c}(Oh)$ and $\theta_{c}(Oh)$, extending this double-threshold framework to a broader class of impurity-triggered, geometry-sensitive instabilities.

### v4

@@ Abstract
In this Letter, we show that an air bubble entrained in a draining liquid sheet can nucleate a hole long before molecular-scale forces become relevant. Axisymmetric direct numerical simulations performed with the Navier-Stokes and combined level-set-volume-of-fluid solver Basilisk C [25, 26] isolate this impurity-driven pathway under controlled conditions while retaining realistic rim dynamics. To focus on the essential physics, we replace the inflating bag geometry of air-blast breakup by an initially flat sheet of thickness $h_{0}$ that drains radially with $u(r)=\omega r$ (fig. 1b), resulting in sheet-thickness far away from the bubble as $h=h_{0}e^{-2\omega t}$, a template that captures both external-airflow forcing in bag breakup [38] and inertia-driven drainage after drop impact [39]. The dynamics are governed by the Ohnesorge number $Oh=\eta/\sqrt{\rho\gamma R_{0}}$, the Bond number $Bo=\rho\omega^{2}R_{0}^{3}/\gamma$, and the non-dimensionlized offset $\chi/R_{0}$ of the bubble’s center from the centerlines of the flow (fig. 1b). Here $\eta$ and $\rho$ are the liquid viscosity and density, $\gamma$ is the surface tension, and $R_{0}$ the bubble radius. The driving acceleration $\omega^{2}R_{0}$ acts as an effective radial gravity. In addition, the initial cavity distortion is characterized by an open angle $\theta$ (fig. 1c), see supplementary material [26] for details on drainage mechanism leading to this geometry.
@@
Our two key results are as follows: First, the entrained bubble grows a through-cavity (a cavity that spans the entire film thickness, forming a continuous opening across the sheet) that overcomes surface tension and nucleates a hole at micrometer-scale thicknesses, explaining rupture without invoking nanoscale physics. Second, the post-nucleation fate is deterministic and requires crossing a double-threshold for irreversible opening: the outward drainage must exceed a critical $Bo_{c}(Oh)$ and the initial cavity distortion – characterized by an opening angle $\theta$ – must exceed a critical $\theta_{c}(Oh)$. Otherwise, rims collide and the sheet heals. The opening-healing boundary collapses onto simple asymptotic scalings: $Bo_{c}=\mathcal{O}(1)$ for $Oh\ll 1$ and $Bo_{c}\sim Oh^{-2}$ for $Oh\gg 1$. These results rationalize the observed perforation of micron-thick sheets by internal defects and provide a predictive framework for impurity-induced breakup in natural and engineered fragmentation flows.
@@
Phenomenology – We first examine sheet-centered bubbles ($\chi/R_{0}=0$) across the $(Oh,Bo)$ space. As the sheet thins, its top and bottom interfaces meet at the bubble poles and a through-cavity opens. In the numerical simulations, the film between the bubble and sheet interface ruptures when the local thickness reaches the smallest grid size, chosen to correspond to approximately 10 nm in physical units to mimic van der Waals-driven rupture. Capillary waves invert the cavity from convex to concave, after which two outcomes are observed (fig. 2a,b): either the rims retract and the hole expands, breaking the sheet, or capillarity drives rim collision and the sheet heals. A systematic sweep in $(Oh,Bo)$ reveals a sharp transition curve (fig. 2c): strong radial driving (large $Bo$) yields opening, whereas weaker driving permits healing unless $Bo$ exceeds a viscosity-dependent threshold. The location and slope of this boundary reflect the competition between capillary-driven rim closure and drainage-driven advection, anticipating the timescale analysis that follows. We additionally show that delayed nucleation–representing chemical/thermal heterogeneity–can be encoded by a larger initial cavity characterized by an opening angle $\theta$, which shifts the transition and thereby introduces geometry as a second way of control.
@@
Figure 2: As the liquid sheet drains radially and thins axially, the interfaces of the bubble and sheet merge to create a cavity. The time instants are shown for $Oh=0.1$, (a) and at small $Bo=10^{-3}$ where the sheet heals, while (b) at larger $Bo=10^{-2}$, the sheet opens up. The left panel depicts velocity magnitude $\|\bm{v}\|/V_{0}$, where $V_{0}=\sqrt{Bo\gamma/\rho R_{0}}$, and the black arrows depict the velocity direction. The right panel illustrates viscous dissipation $\xi=2Bo\left(\bm{\mathcal{D}:\mathcal{D}}\right)$, where $\bm{\mathcal{D}}=\left(\bm{\nabla u+\nabla u}^{T}\right)/2$ is the symmetric part of velocity gradient tensor. (c) Effect of driving and viscosity on sheet rupture. The regime map in the log-log parameter space of OhBoOh-Bo. The transition lines at small OhOh are shown by a constant BoBo line in gray, while, at large OhOh, the transition is indicated by a black line with scaling $Oh\sim Bo^{-1/2}$.
@@
What happens in the limiting case of no external driving ($Bo=0$)? – In the absence of radial forcing ($Bo=0$), the liquid sheet inevitably heals for any finite $Oh$ – even extremely viscous sheets (large $Oh$) will eventually reseal, albeit very slowly as viscosity prolongs the rim collision time. This follows from our thresholds: as $Bo\to 0$, achieving opening would require $Oh\sim Bo^{-1/2}\to\infty$, so any finite $Oh$ lies in the healing regime. Our simulations confirm this universal healing at $Bo=0$, consistent with the classic geometric energy criterion of Taylor & Michael [40]. Using a toroidal model for the hole’s rim [41], they predicted that without external driving a circular hole can continue to expand only if its radius exceeds a purely geometric threshold. In particular, when the hole’s outer radius $R$ [46] satisfies $R/h_{0}>\pi/4$ (so that the inner radius $R_{h}=R-h_{0}/2$ is about $0.29h_{0}$), the excess surface energy $\Delta\Xi$ becomes positive, signaling that the hole will grow rather than close. Holes smaller than this critical size have $\Delta\Xi<0$ and will heal. Notably, this threshold is independent of viscosity. Thus, at $Bo=0$ no finite-$Oh$ case can open unless the initial hole is above the critical $R/h_{0}$ – indeed, the cavities in our simulations were below this size, so they all heal.
In more realistic situations, additional physics can delay the initial rupture (for example, due to chemical or thermal heterogeneity), allowing the sheet to drain longer and thin more broadly before the hole forms. We represent such delays in our model by a larger initial cavity distortion characterized by an increased polar angle $\theta$ (fig. 1c). Simulations at $Bo=0$ with varied $\theta$ reveal a clear geometry-controlled threshold: for sufficiently large distortions (large $\theta$), the hole opens even without any driving, whereas for smaller initial distortions the hole heals (fig. 3b). This confirms the second threshold of the double-threshold framework: a sufficiently large initial cavity alone can trigger sheet rupture in the absence of external forcing. Moreover, at very high viscosity (large Oh), the opening–healing transition occurs at a nearly constant opening angle $\theta\approx 0.09\pi$. Remarkably, this transition, when expressed in terms of $R_{h}/h_{0}$, gives the same result as the Taylor–Michael criterion $R_{h}/h_{0}\approx 0.29$. However, as $Oh$ decreases (inertia-dominated regime), this critical distortion $\theta_{c}$ rises, since added inertia aids rim closure—requiring a larger initial opening to overcome healing. Remarkably, sheets that do open exhibit non-monotonic cavity growth (fig. 3c). Surface tension first drives the cavity edges apart, causing radial cavity growth, but the induced inertia pulls the rim inward and temporarily reduces the cavity radius. However, the cavity remains large enough to overcome healing by this recoil and eventually opens with the Taylor–Culick mechanism.
@@
Figure 4: The draining liquid sheet of the bubble is placed asymmetrically with $\chi/R_{0}=0.1$. Time evolution has been shown for cases with $Oh=1$, (a) $Bo=10^{-4}$, and (b) $Bo=10^{-2}$. In the former case, the liquid bridge at the south pole replenishes and grows into a flat-shaped sheet due to capillary action, and then the sheet keeps on thinning due to radial drainage. Meanwhile, in the latter case, thinning due to radial drainage dominates before the bridge can be replenished by the damped capillary waves at large $Oh=1$. The left panel depicts velocity magnitude $\|\bm{v}\|/V_{0}$, where $V_{0}=\sqrt{Bo\gamma/\rho R_{0}}$, the black arrows depict the velocity direction. The right panel depicts viscous dissipation $\xi=2Bo\left(\bm{\mathcal{D}:\mathcal{D}}\right)$, where $\bm{\mathcal{D}}=\left(\bm{\nabla u+\nabla u}^{T}\right)/2$ is the symmetric part of velocity gradient tensor. (c) Effect of the asymmetric bubble position on sheet breakup. Sheet breakup time (normalized by the no-bubble case rupture time $\tau^{*}=3.5/\omega\equiv 3.5\tau_{\gamma}/\sqrt{Bo}$, with the prefactor set by the grid cutoff) in the OhBoOh-Bo parameter space for $\chi/R_{0}=0.1$. The plot is constructed from simulation results at 324 points arranged on a uniform logarithmic $18\times 18$ grid in parameter space. The gray line depicts the transition line observed for symmetric cases ($\chi/R_{0}=0$), as shown in fig. 2.
@@
A trapped bubble – acting as a hydrophobic defect – thus rationalizes why rupture occurs far above molecular scales and provides predictive control knobs, namely driving strength and defect geometry, relevant to bag-breakup sprays and respiratory films. When the bubble is off-center, however, the sharp $Oh_{c}\sim Bo^{-1/2}$ transition gives way to a gradual crossover governed by the competing timescales of capillary-wave replenishment, viscous damping, and radial drainage – physics that awaits a predictive scaling theory. In practice, small off-center bubbles at low $Oh$ and $Bo$ are unlikely to trigger rupture unless delayed nucleation (e.g., from chemical or thermal heterogeneity) enlarges the initial cavity sufficiently. Future studies could test whether analogous thresholds govern rupture triggered by solid particles, oil droplets, or Marangoni-driven inhomogeneities. Chemical or thermal gradients may couple nonlinearly with radial drainage and shift $Bo_{c}(Oh)$ and $\theta_{c}(Oh)$, extending this double-threshold framework to a broader class of impurity-triggered, geometry-sensitive instabilities.